# Strategic Content Creation in the Age of GenAI: To Share or Not to Share?

### v1

@@ C.2 Complete Proof of Theorem 6
Therefore, by taking $B=n\cdot(1+L_{0}\cdot\max_{i}x_{i}^{\max})$, we can express the time complexity as $O\left(\frac{B}{\delta^{2}}\right)=O\left(\frac{A^{2}B}{\varepsilon^{2}}\right)$, where $A$ is the constant we defined earlier. This completes the proof of the computational complexity and ends the proof of Theorem 6. ∎
Remark 24.
Consider an instance $\mathcal{I}$ with $n$ creators, where each creator $i$ has a cost function $c_{i}$ satisfying the following regularity conditions:
  1. 1.
    There exist constants $a_{1},\theta_{1},a_{2},\theta_{2}>0$ with $1<\theta_{1}\leq\theta_{2}$ such that
    $a_{1}x^{\theta_{1}}\leq c_{i}(x)\leq a_{2}x^{\theta_{2}}\quad\text{for all }x\geq 1.$
  2. 2.
    The inverse of the derivative $(c_{i}^{\prime})^{-1}$ exists and satisfies, for all $y\geq y_{0}$,
    $K_{1}y^{1/\psi_{1}}\leq(c_{i}^{\prime})^{-1}(y)\leq K_{2}y^{1/\psi_{2}}.$
    for some $K_{1},K_{2},y_{0}>0$ and $0<\psi_{1}\leq\psi_{2}$.
Then the algorithmic complexity constants $A=A(\mathcal{I})$ and $B=B(\mathcal{I})$ from Theorem 6 are polynomial in $n$.
Proof of Remark 24.
We explicitly show that each quantity entering $A$ and $B$ is polynomial in $n$ under the stated regularity conditions, by expressing them in terms of the upper and lower bounds, and verifying these bounds are polynomial in $n$.
Step 1: Upper and lower bounds on $x_{i}^{\max}$, $X_{\mathrm{UB}}$, $Y_{\mathrm{UB}}$.
By the properties of $c_{i}$, for each $i$, $x_{i}^{\max}$ (the solution to $\mu(x_{i}^{\max})^{\gamma}=c_{i}(x_{i}^{\max})$) is bounded above and below by constants independent of $n$. Therefore,
$X_{\mathrm{UB}}=\sum_{i=1}^{n}x_{i}^{\max}=\Theta(n),\quad Y_{\mathrm{UB}}=X_{\mathrm{UB}}-\min_{k}x_{k}^{\max}=\Theta(n).$
Step 2: Lower bounds $X_{\mathrm{LB}}$, $Y_{\mathrm{LB}}$.
Similarly, $X_{\mathrm{LB}}=\sum_{i=1}^{n}(c_{i}^{\prime})^{-1}(y)$ for some $y$ polynomial in $n$ (since $y=\mu\gamma(X_{\mathrm{UB}})^{\gamma-1}$ and $X_{\mathrm{UB}}=\Theta(n)$). By regularity, $(c_{i}^{\prime})^{-1}(y)$ is polynomial in $n$ for each $i$, so
$X_{\mathrm{LB}}=\Theta(n^{a})\text{ and }Y_{\mathrm{LB}}=\Theta(n^{b}).$
for some exponents $a,b>0$, i.e., both are polynomial in $n$.
Step 3: The constants $L$, $M_{1}$, $M_{2}$.
Each of these constants is defined directly in terms of the upper and lower bounds above:
  • The Lipschitz constant
    $L=\frac{\mu\sqrt{n}}{\beta}\frac{X_{\mathrm{UB}}}{(X_{\mathrm{LB}})^{2-\gamma}},$
    which is a rational function of $n$, $X_{\mathrm{UB}}$, $X_{\mathrm{LB}}$ (each polynomial in $n$), and fixed model parameters, and thus itself polynomial in $n$.
  • The constant
    $M_{1}=M_{c^{\prime}}+\mu\cdot B(0.5Y_{\mathrm{LB}},1.5Y_{\mathrm{UB}},n),$
    where $M_{c^{\prime}}$ is the maximum derivative of $c_{i}$ over $[0,x_{i}^{\max}]$ (bounded by a polynomial in $n$), and $B(Y_{\mathrm{LB}},Y_{\mathrm{UB}},n)$ is a function of $Y_{\mathrm{LB}}$, $Y_{\mathrm{UB}}$, and $n$ with only addition, multiplication, and exponentiation by constants, so it is polynomial in $n$.
  • Similarly,
    $M_{2}=\mu\left(\gamma(0.5Y_{\mathrm{LB}})^{\gamma-1}\sqrt{n}+(1.5Y_{\mathrm{UB}})^{\gamma}\right),$
    which is a sum and product of polynomials in $n$ and fixed parameters, so also polynomial in $n$.
Therefore, $L$, $M_{1}$, and $M_{2}$ are polynomial in $n$ as they are built from polynomially bounded quantities.
Step 4: Complexity constants $A$ and $B$.
Recall that
[…]
and
$B=n\cdot(1+L\cdot\max_{i}x_{i}^{\max}).$
Since $L$, $M_{1}$, $M_{2}$, $X_{\mathrm{UB}}$, $Y_{\mathrm{LB}}$, and $\max_{i}x_{i}^{\max}$ are all polynomial in $n$, their sums, products, and maxima are also polynomial in $n$. Thus, $A$ and $B$ are polynomial in $n$. ∎
Appendix D Omitted Simulations Details
This section provides technical details about the algorithms and computational resources used in our experimental simulations.
D.1 Multi-Agent Mirror Descent
Algorithm 3 approximates the unique ESE for a given revenue allocation parameter $\rho$ (always exists due to Proposition 3 and guaranteed to converge due to the game being strongly monotone (Proposition 15)). It uses gradient-based updates to iteratively refine each creator’s quality choice $x_{i}$ while maintaining non-negativity constraints.
Algorithm 3 Multi-Agent Mirror Descent for ESE Quality Profile ($x^{\star}$)
2:Approximate ESE quality profile $x^{(T)}$.
3:Initialize quality profile $x\leftarrow x^{(0)}$.
4:for $t=0$ to $T-1$ do
5:     Calculate gradient vector $v^{(t)}\in\mathbb{R}^{n}$. For each creator $i\in\{1,\dots,n\}$:
6:     Let $X^{(t)}=\sum_{j=1}^{n}x_{j}^{(t)}$ and $X_{-i}^{(t)}=X^{(t)}-x_{i}^{(t)}$.
7:     Set $K(\rho)=\frac{\mu(1+\alpha\rho)}{1+\alpha}$.
8:     $v_{i}^{(t)}\leftarrow K(\rho)(X^{(t)})^{\gamma-2}[\gamma x_{i}^{(t)}+X_{-i}^{(t)}]-c_{i}^{\prime}(x_{i}^{(t)})$. $\triangleright$ Utility Gradient
9:     Update candidate profile: $x_{cand}\leftarrow x^{(t)}+\eta_{t}v^{(t)}$.
10:     Project onto feasible set: $x^{(t+1)}\leftarrow\max(0,x_{cand})$. $\triangleright$ Component-wise max
11:return $x^{(T)}$.
D.2 FSE Stability Check
Algorithm 4 verifies whether an FSE by checking if any creator can profitably deviate to not sharing ($s_{i}=0$) (necessary and sufficient due to Proposition 2). For each creator, it solves a one-dimensional optimization problem to find their best response when withholding content. The profile is deemed stable only if no creator benefits from such deviation beyond the tolerance threshold $\varepsilon_{tol}$. Notice that it is similar to Algorithm 2, but in the empirical simulations, we used a standard 1D solver rather than a grid search.
Algorithm 4 FSE Stability Check via $s_{i}=0$ Deviation
3:for $i=1$ to $n$ do
4:     Calculate utility at candidate profile: $U_{i}^{\star}\leftarrow U_{i}(x_{i}^{\star},1;x_{-i}^{\star},\mathbf{1}_{-i})$.
5:     Calculate rational quality upper bound for deviation: $x_{i}^{max}\leftarrow(\mu/a_{i})^{1/(\theta_{i}-\gamma)}$.
6:     $U_{dev}(s_{i}=0)\leftarrow\max_{x_{i}^{\prime}\in[0,x_{i}^{max}]}\{U_{i}(x_{i}^{\prime},0;x_{-i}^{\star},\mathbf{1}_{-i})\}$. $\triangleright$ Requires 1D solver
7:     Check stability condition:
8:     if $U_{dev}(s_{i}=0)-U_{i}^{\star}>\varepsilon_{tol}$ then
9:         return False $\triangleright$ Player $i$ has a profitable deviation
D.3 Simulation Pipeline
For each simulation run:
  1. 1.
    For each parameter combination we checked, we initialize 150 independent game instances with randomly sampled cost coefficients.
  2. 2.
    For each instance, we evaluate 100 equally-spaced values of $\rho\in[0,1]$.
  3. 3.
    Algorithm 3 computes the ESE for each $\rho$ value.
  4. 4.
    Algorithm 4 verifies whether each ESE is also an FSE.
  5. 5.
    Results are averaged across instances with 95% confidence intervals computed via bootstrapping.
D.4 Computational Resources and Runtime
The complete simulation suite was executed on a standard MacBook Pro with 18 GB of RAM and 8 CPU cores. We utilized parallel processing across 6 cores to accelerate computation. The full set of experiments required approximately 6 hours of computation time. The breakdown includes:
  • ESE computation and FSE verification across all parameter combinations. This took about 6 hours in total.
  • Statistical analysis and bootstrapping for confidence intervals, plotting and visualization. Took less than 1 minute.
The implementation uses standard Python scientific computing libraries (NumPy, SciPy) without GPU acceleration, as the algorithms are primarily CPU-bound.
Appendix E FSE Characterization Under Power Cost Functions
In this Section, we further continue the analysis of FSE existence and platform optimization by focusing on a family of cost functions with strong structural properties. Specifically, we consider the class of power cost functions, which both satisfy our general assumptions and allow us to derive sharper insights.
Definition 3.
An $(n,a_{\min},a_{\max},\theta)$-power cost instance is an instance consisting of $n$ creators, and for each $i\in[n]$, the cost function takes the form $c_{i}(x_{i})=a_{i}x_{i}^{\theta}$, where $\theta\in(1,2]$ is a common exponent, and the cost coefficients satisfy $a_{i}\in[a_{\min},a_{\max}]$.
Power cost functions are standard in economic and game-theoretic models of effort and production (e.g., [54, 38, 52]). They offer a tractable yet flexible structure for modeling convex production costs.
E.1 Asymptotic FSE Existence in Large Ecosystems
One of the main technical challenges in FSE characterization is that the stability condition derived in Theorem 5 depends on the equilibrium outcome itself. While exact evaluation is intractable, power cost functions enable asymptotic reasoning. In particular, we show that as the number of creators grows large, inducing full sharing becomes increasingly easier to induce, even under modest revenue-sharing rates.
Proposition 25.
Assume a large enough $\mu$. For any $n,a_{\min},a_{\max},\theta$ and $w\in(0,1]$, there exists a threshold $N(w)$ such that for all $n>N(w)$, any $(n,a_{\min},a_{\max},\theta)$-power cost instance with $n$ creators admits an FSE for all revenue-sharing rates $\rho\in(w,1]$.
This result shows that the platform need not share large portions of GenAI-driven revenue to achieve full sharing in large-scale creator ecosystems. The intuition is that as the number of creators increases, the total content pool grows, which reduces the marginal incentive to deviate from sharing (see discussion around thresholds in Proposition 2).
Proof of Proposition 25.
We aim to determine an integer threshold $N(w)$ such that for any game with $n>N(w)$ creators, the condition from Theorem 5 holds for all $\rho\in(w,1]$. This condition ensures that the unique ESE $\mathcal{X}(\rho)$ also constitutes the unique FSE $(\mathcal{X}(\rho),\mathbf{1})$, and is given by:
$\rho>\max_{i\in[n]}\left\{\frac{x_{i}^{\max}}{x_{i}^{\max}+(1+\alpha)\left\|\mathcal{X}(\rho)_{-i}\right\|}\right\}\eqqcolon\max_{i\in[n]}\tau_{i}(\rho).$ (4)
We establish this by showing that the RHS of Equation 4 vanishes as $n\to\infty$. Furthermore, we also provide a threshold $N(w)$ for this condition to hold with respect to all values of $\rho$ such that $\rho\geq w$. For ease of notation, we will use the notation $\tau_{i}(\rho)\coloneqq\frac{x_{i}^{\max}}{x_{i}^{\max}+(1+\alpha)\left\|\mathcal{X}(\rho)_{-i}\right\|}$ for the rest of the proof.
Step 1: asymptotic lower bound for $\left\|\mathcal{X}(\rho)_{-i}\right\|$.
We first analyze the behavior of the equilibrium qualities in the enforced sharing game as the number of creators $n$ increases.
To that end, we use a reduction from the enforced sharing game to the exclusive competition model of Yao et al. [52]. Under enforced sharing, the utility of player $i$ is given by Lemma 9. It turns out that the 1D exclusive competition in Yao et al. [52] is equivalent to our enforced sharing game with the following parameter substitution:$\mu^{\prime}=\mu(1+\alpha\rho)$, $\tilde{\gamma}=\gamma$, $\tilde{\beta}=1$ and $\tilde{\alpha}=\alpha$.
Thus, we can apply their results about power costs in order to tightly characterize how the total human content quality behaves asymptotically. Specifically, we can adapt Theorem 3 and Lemma 6 from Yao et al. [52] to our use case. The adapted versions of the results are written below. It is important to note that our main results in this paper are novel and require new ideas and techniques beyond those covered in Yao et al. [52]. More precisely, our two main points of interest were the sharing decisions of the creators and the allocation rule of the platform, which are novel aspects of our paper.
Theorem 26 (Adapted from Theorem 3 in [52]).
Assume power cost functions $c_{i}(x_{i})=a_{i}x_{i}^{\theta}$ for all $i$. Fix any $\rho\in[0,1]$ and let $\mu^{\prime}=\mu(1+\alpha\rho)$. For any sufficiently large $n$ and $\mu$, $\left\|\mathcal{X}(\rho)\right\|$ satisfies:
$\frac{C_{\theta}}{2\alpha+2}<\frac{(\left\|\mathbf{x}^{\star}\right\|)^{\gamma+\theta-1}}{\mu^{\prime}\cdot\left\|\mathbf{a}^{-1}\right\|_{\frac{1}{\theta-1}}^{\theta-1}}<C_{\theta},$
where $C_{\theta}$ is a constant depending only on $\theta$, $\mathbf{a}^{-1}=(a_{1}^{-1},\dots,a_{n}^{-1})$, and $\left\|\mathbf{a}^{-1}\right\|_{\frac{1}{\theta-1}}=\left(\sum_{i=1}^{n}(a_{i}^{-1})^{\frac{1}{\theta-1}}\right)$.
Lemma 27 (Adapted from Lemma 6 in [52]).
Assume power cost functions $c_{i}(x_{i})=a_{i}x_{i}^{\theta}$ for all $i$. Fix any $\rho\in[0,1]$ and let $\mu^{\prime}=\mu(1+\alpha\rho)$. For any sufficiently large $n$ and $\mu$,
$\max_{i\in\{1,\dots,n\}}\{\mathcal{X}(\rho)_{i}\}\leq\frac{\left\|\mathcal{X}(\rho)\right\|}{2}.$
Theorem 26 suggests there exists a constant $N_{YN}$ (a threshold on $n$) such that for any game with $n>N_{YN}$ creators:
$(\left\|\mathcal{X}(\rho)\right\|)^{\gamma+\theta-1}>\frac{C_{\theta}}{2\alpha+2}\mu^{\prime}\left(\sum_{j=1}^{n}a_{j}^{-1/(\theta-1)}\right)^{\theta-1}.$
To get a uniform lower bound for $\rho\in[w,1]$, we use $\mu^{\prime}=\mu(1+\alpha\rho)\geq\mu(1+\alpha w)$. We also bound the sum using the maximum cost coefficient $a_{\max}$ (which implies $a_{j}^{-1/(\theta-1)}\geq a_{\max}^{-1/(\theta-1)}$): $\sum_{j=1}^{n}a_{j}^{-1/(\theta-1)}\geq na_{\max}^{-1/(\theta-1)}$. Substituting these bounds yields the lower bound on total quality for $n>N_{YN}$ and $\rho\in[w,1]$:
$\left\|\mathcal{X}(\rho)\right\|>\left[\frac{C_{\theta}}{2\alpha+2}\mu(1+\alpha w)(na_{\max}^{-(\theta-1)})^{\theta-1}\right]^{1/(\gamma+\theta-1)}\eqqcolon LB_{n}(w).$
This shows $\left\|\mathcal{X}(\rho)\right\|$ grows unbounded as $n\to\infty$, lower-bounded by $LB_{n}(w)$.
Furthermore, Lemma 27 states that for all $n>N_{YN}$ (and $\theta\in(1,2]$), the maximum individual quality is bounded relative to the total quality: $\max_{j\in[n]}\mathcal{X}(\rho)_{j}\leq\frac{\left\|\mathcal{X}(\rho)\right\|}{2}$. This allows us to lower-bound the sum of qualities of creators other than $i$:
$\left\|\mathcal{X}(\rho)_{-i}\right\|=\left\|\mathcal{X}(\rho)\right\|-\mathcal{X}(\rho)_{i}\geq\left\|\mathcal{X}(\rho)_{-i}\right\|-\max_{j\in[n]}\mathcal{X}(\rho)_{j}\geq\left\|\mathcal{X}(\rho)\right\|-\frac{\left\|\mathcal{X}(\rho)\right\|}{2}=\frac{\left\|\mathcal{X}(\rho)\right\|}{2}.$
Combining with the lower bound on $\left\|\mathcal{X}(\rho)\right\|$, we get the crucial asymptotic lower bound for $\left\|\mathcal{X}(\rho)_{-i}\right\|$ for $n>N_{YN}$ and $\rho\in[w,1]$:
$\left\|\mathcal{X}(\rho)_{-i}\right\|\geq\frac{1}{2}LB_{n}(w).$ (5)
Step 2: bounding the threshold and deriving $N(w)$.
We need to ensure that for $n>N(w)$, condition (4) holds for all $\rho\in(w,1]$. Since $\rho>w$, it is sufficient to guarantee that the maximum threshold term is less than or equal to $w$, i.e., $\max_{i}\tau_{i}(\rho)\leq w$. First, we find a uniform upper bound for $x_{i}^{\max}$. Recall $x_{i}^{\max}$ is the solution to $a_{i}(x_{i}^{\max})^{\theta}=\mu(x_{i}^{\max})^{\gamma}$, which gives:
$x_{i}^{\max}=\left(\frac{\mu}{a_{i}}\right)^{\frac{1}{\theta-\gamma}}\leq\left(\frac{\mu}{a_{\min}}\right)^{\frac{1}{\theta-\gamma}}\eqqcolon M_{qual},$
where $M_{qual}$ is a finite positive constant independent of $n$ and $\rho$. Now we bound $\tau_{i}(\rho)$ using this and the lower bound for $\left\|\mathcal{X}(\rho)\right\|$ from Step 1, for $n>N_{YN}$ and $\rho\in[w,1]$:
$\tau_{i}(\rho)=\frac{x_{i}^{\max}}{x_{i}^{\max}+(1+\alpha)\left\|\mathcal{X}(\rho)_{-i}\right\|}\leq\frac{M_{qual}}{M_{qual}+(1+\alpha)\frac{1}{2}LB_{n}(w)}.$
To ensure $\max_{i}\tau_{i}(\rho)\leq w$, it suffices to impose this condition on the upper bound:
$\frac{M_{qual}}{M_{qual}+(1+\alpha)\frac{1}{2}LB_{n}(w)}\leq w.$
We rearrange the inequality:
$M_{qual}\leq wM_{qual}+w(1+\alpha)\frac{1}{2}LB_{n}(w),$
$(1-w)M_{qual}\leq\frac{w(1+\alpha)}{2}LB_{n}(w),$
$LB_{n}(w)\geq\frac{2(1-w)M_{qual}}{w(1+\alpha)}.$
Substitute the expressions for $LB_{n}(w)$ and $M_{qual}$:
$\left[\frac{C_{\theta}}{2\alpha+2}\mu(1+\alpha w)(na_{\max}^{-(\theta-1)})^{\theta-1}\right]^{1/(\gamma+\theta-1)}\geq\frac{2(1-w)}{w(1+\alpha)}\left(\frac{\mu}{a_{\min}}\right)^{\frac{1}{\theta-\gamma}}.$
Raise both sides to the power $\gamma+\theta-1>0$:
$\frac{C_{\theta}}{2\alpha+2}\mu(1+\alpha w)n^{\theta-1}a_{\max}^{-(\theta-1)}\geq\left[\frac{2(1-w)}{w(1+\alpha)}\left(\frac{\mu}{a_{\min}}\right)^{\frac{1}{\theta-\gamma}}\right]^{\gamma+\theta-1}.$
Isolate $n^{\theta-1}$:
$n^{\theta-1}\geq\frac{(2\alpha+2)a_{\max}^{\theta-1}}{C_{\theta}\mu(1+\alpha w)}\left[\frac{2(1-w)}{w(1+\alpha)}\left(\frac{\mu}{a_{\min}}\right)^{\frac{1}{\theta-\gamma}}\right]^{\gamma+\theta-1}.$ (6)
Let $RHS(w)$ denote the RHS of Equation 6. The inequality $n^{\theta-1}\geq RHS(w)$ must hold. The smallest integer $n$ for which this holds is given by $\lceil(RHS(w))^{1/(\theta-1)}\rceil$ (if $RHS(w)>0$, otherwise $n=1$). Let $N_{bound}(w)$ be this value:
$N_{bound}(w)=\left\lceil\left\{\frac{(2\alpha+2)a_{\max}^{\theta-1}}{C_{\theta}\mu(1+\alpha w)}\left[\frac{2(1-w)}{w(1+\alpha)}\left(\frac{\mu}{a_{\min}}\right)^{\frac{1}{\theta-\gamma}}\right]^{\gamma+\theta-1}\right\}^{\frac{1}{\theta-1}}\right\rceil.$
Choosing the overall threshold as $N(w)=\max(N_{YN},N_{bound}(w))$ guarantees that for all games with $n>N(w)$ creators, the condition $\max_{i}\tau_{i}(\rho)\leq w$ holds for all $\rho\in[w,1]$. Since we consider $\rho\in(w,1]$, we have $\rho>w\geq\max_{i}\tau_{i}(\rho)$, satisfying condition (4). Therefore, by Theorem 5, for $n>N(w)$ and $\rho\in(w,1]$, the qualities profile of the unique ESE $\mathcal{X}(\rho)$ also constitutes an FSE $(\mathcal{X}(\rho),\mathbf{1})$. ∎
E.2 Monotonicity of Content Production w.r.t. $\rho$
A natural question is how the platform’s revenue-sharing decision affects total content quality. We answer this in the positive by showing that when costs are symmetric power functions, content production increases monotonically with $\rho$. This offers important design guidance: Even if full sharing can be achieved with low $\rho$, choosing a higher value may yield more human content-and hence more user traffic.
Proposition 28.
Fix any instance with symmetric power costs, i.e., $a_{i}=a$ for all $i\in[n]$ and some $a>0$. Then $\left\|\mathcal{X}(\rho)\right\|$ is strictly increasing in $\rho$.
This result highlights a key trade-off: Increasing $\rho$ reduces the platform’s retained share of AI-driven revenue, but incentivizes higher creator effort, which increases overall engagement and traffic.
Proof of Proposition 28.
Under the given assumptions, the enforced-sharing game is symmetric, with homogeneous cost functions $c_{i}(x)=ax^{\theta}$ for all $i$, where $a>0$ and $\theta>1$. Proposition 3 guarantees a unique symmetric ESE with $\mathcal{X}(\rho)_{i}=x(\rho)$ for all $i$.
We first invoke Lemma 12, which ensures that any rational equilibrium quality must be strictly positive. Therefore, $x(\rho)>0$ for all $\rho\in[0,1]$, and the total quality is $\left\|\mathcal{X}(\rho)\right\|=n\cdot x(\rho)$ with $n\geq 2$.
To analyze how $\mathcal{X}(\rho)$ varies with $\rho$, we use the first-order condition (FOC) characterizing equilibrium under full sharing. Since $\mathcal{X}(\rho)>0$, the FOC is valid and given by the comparing the derivation of the utility (as presented in Lemma 9) with respect to $x_{i}$ to 00:
$K(\rho)\cdot(\left\|\mathcal{X}(\rho)\right\|)^{\gamma-2}\left[(\gamma-1)x(\rho)+\left\|\mathcal{X}(\rho)\right\|\right]=c^{\prime}(x(\rho)),$
where $K(\rho)=\frac{\mu(1+\alpha\rho)}{1+\alpha}$. Substituting $\left\|\mathcal{X}(\rho)\right\|=nx(\rho)$ and $c^{\prime}(x(\rho))=a\theta(x(\rho))^{\theta-1}$, we obtain:
$K(\rho)\cdot(nx(\rho))^{\gamma-2}x(\rho)(n+\gamma-1)=a\theta(x(\rho))^{\theta-1}.$
Dividing both sides by $x(\rho)^{\gamma-1}$ (which is positive), we simplify:
$K(\rho)\cdot n^{\gamma-2}(n+\gamma-1)=a\theta(x(\rho))^{\theta-\gamma}.$
Define $A\coloneqq\frac{n^{\gamma-2}(n+\gamma-1)}{a\theta}>0$, a constant independent of $\rho$. Then,
$x(\rho)=\left[A\cdot K(\rho)\right]^{\frac{1}{\theta-\gamma}}.$
Since $\theta>\gamma$ (because $\theta>1\geq\gamma$), the exponent $\frac{1}{\theta-\gamma}$ is positive, and $x(\rho)$ is strictly increasing in $K(\rho)$.
Now consider the derivative of $K(\rho)$:
$\frac{dK(\rho)}{d\rho}=\frac{\mu\alpha}{1+\alpha},$
which is strictly positive whenever $\alpha>0$. Therefore, $K(\rho)$ is strictly increasing in $\rho$, which implies that $x(\rho)$ is strictly increasing in $\rho$ as well.
Finally, since $\left\|\mathcal{X}(\rho)\right\|=n\cdot x(\rho)$, the total equilibrium quality is also strictly increasing in $\rho$. ∎
E.3 Constant-Factor Approximation via Closed-Form Allocation Parameter
Finally, we derive a closed-form choice of $\rho$ that yields a provable approximation to the platform’s optimal revenue. This result provides a simple, interpretable heuristic that is especially useful when solving the full optimization problem is computationally infeasible.
Proposition 29.
Assume large enough $n,\mu$, and that all creators have power costs $c_{j}(x)=a_{j}x^{\theta}$ with $a_{j}>0$ and common exponent $\theta\in(1,2]$, such that $\alpha\gamma+1>\gamma+\theta$. Then, setting
$\rho=\frac{\alpha\gamma-\gamma-\theta+1}{\alpha(2\gamma+\theta-1)}$
ensures that $\rho$ induces an FSE, and yields platform revenue satisfying
$U_{p}(\mathcal{X}(\rho),\mathbf{1};f_{\rho})\geq(2\alpha+2)^{-\frac{\gamma}{\gamma+\theta-1}}\cdot\text{OPT}.$
The expression for $\rho$ depends only on structural parameters of the model and is independent of any details on the creators, besides the assumption that they have power costs. This result implies that platforms can use a fixed rule, without solving a bi-level optimization, to obtain provably good outcomes, particularly in large systems.
Proof of Proposition 29.
We begin by reestablishing key relationships between $\left\|\mathcal{X}(\rho)\right\|$ and its lower bound $X_{LB}(\rho)$. Theorem 26 guarantees that for any $\rho\in\mathcal{F}$:
$X_{LB}(\rho)<\left\|\mathcal{X}(\rho)\right\|<R\cdot X_{LB}(\rho),$
where $R=(2\alpha+2)^{1/(\gamma+\theta-1)}$ and $X_{LB}(\rho)=\left(\frac{C_{\theta}\mu(1+\alpha\rho)N_{norm}}{2\alpha+2}\right)^{\frac{1}{\gamma+\theta-1}}$.
Since the platform’s revenue is $U_{p}(\mathcal{X}(\rho),\mathbf{1};f_{\rho})=\frac{\mu\alpha}{1+\alpha}(1-\rho)(X^{\star}(\rho))^{\gamma}$, these quality bounds imply corresponding revenue bounds:
$V_{LB}(\rho)<U_{p}(\mathcal{X}(\rho),\mathbf{1};f_{\rho})<R^{\gamma}\cdot V_{LB}(\rho),$
@@ E.3 Constant-Factor Approximation via Closed-Form Allocation Parameter
$V_{LB}(\rho)$ $=\frac{\mu\alpha}{1+\alpha}(1-\rho)\left[\left(\frac{C_{\theta}\mu(1+\alpha\rho)N_{norm}}{2\alpha+2}\right)^{\frac{1}{\gamma+\theta-1}}\right]^{\gamma}$

### v2

@@ C.5 Proofs of Section 4’s Main Results
Therefore, by taking $B=n\cdot(1+B(1,5\cdot 2^{5-\gamma},0,2,3-\gamma)\cdot x^{\max})$, we can express the time complexity as $O\left(\frac{B}{\delta^{2}}\right)=O\left(\frac{A^{2}B}{\varepsilon^{2}}\right)$, where $A$ is the constant we defined earlier. This completes the proof of the computational complexity and ends the proof of Theorem˜8. ∎
See 9
Proof of Proposition˜9.
We prove the proposition by using the tight bounds on the ESE qualities for power cost functions established in Lemma˜16. Notice that this does not alter any of the previously mentioned results, as it still correctly bounds the ESE qualities from above and below.
Step 1: constants simplification
Notice that under the assumption that $a_{\min},a_{\max}=O(1)$, we have $x^{\max},M_{c^{\prime}}=O(1)$. Additionally, from Lemma˜16, we have that $Y_{LB},Y_{UB}=\Theta\left(n^{\frac{\theta-1}{\theta-\gamma}}\right)$. Therefore, we can simplify the function $B(\cdot)$ as follows:
$B(q,u,v,w,t)=q\cdot M_{c^{\prime}}+\frac{un^{v}\left(1+x^{\max}+Y_{UB}\right)^{w}}{(Y_{LB})^{t}}=O\left(n^{\max\{0,v+\frac{\theta-1}{\theta-\gamma}\cdot(w-t)\}}\right).$ (12)
Step 2: Analyzing the constants in Theorem 8
We now analyze the constants that appear in the proof of Theorem 8. Starting with the constant $A$, we have:
$A$ $=4B(1,(5+\alpha)\cdot 2^{6-\gamma},0.5,3,3-\gamma)\cdot(1+B(0,1/\beta,0.5,1,2-\gamma))$
$=O\left(n^{\max\{0,0.5+\frac{\theta-1}{\theta-\gamma}\cdot(3-(3-\gamma))\}}\cdot n^{\max\{0,0.5+\frac{\theta-1}{\theta-\gamma}\cdot(1-(2-\gamma))\}}\right)$
$=O\left(n^{0.5+\frac{\theta-1}{\theta-\gamma}\cdot\gamma+\max\{0,0.5+\frac{\theta-1}{\theta-\gamma}\cdot(\gamma-1)\}}\right).$
In our regime of interest, where $\gamma\in[0,1]$ and $\theta\in(1,2]$, the power of $n$ in the above expression is maximized at $\gamma=1$, and we have
$0.5+\frac{\theta-1}{\theta-\gamma}\cdot\gamma+\max\{0,0.5+\frac{\theta-1}{\theta-\gamma}\cdot(\gamma-1)\}\leq 2.$
Therefore, we have $A=O\left(n^{2}\right)$. Moving on to the constant $B$ from the complexity analysis, we have:
$B$ $=n\cdot(1+B(1,5\cdot 2^{5-\gamma},0,2,3-\gamma)\cdot x^{\max})$
$=O\left(n\cdot n^{\max\{0,0+\frac{\theta-1}{\theta-\gamma}\cdot(\gamma-1)\}}\right)=O(n).$
To conclude, we have $A=O\left(n^{2}\right)$ and $B=O(n)$, which completes the proof of Proposition˜9. ∎
See 10
Proof of Proposition˜10.
We begin by invoking a result adapted from Yao et al. [50]. We use a reduction from the enforced sharing game to the exclusive competition model analyzed in Yao et al. [50]. Under enforced sharing, the utility of player $i$ is given by Lemma˜12. It turns out that the 1D exclusive competition in Yao et al. [50] is equivalent to our enforced sharing game with the following parameter substitution: $\mu^{\prime}=\mu(1+\alpha\rho)$, $\tilde{\gamma}=1-\gamma$, $\tilde{\beta}=1$ and $\tilde{\alpha}=\alpha$.
Yao et al. [50] derived asymptotic bounds on the total quality in the PNE of the exclusive competition, which directly translate to bounds for our ESE concept. Specifically, we adapt Theorem 3 from Yao et al. [50] to our use case. We stress again that our main results in this paper are novel and require new ideas and techniques beyond those covered in Yao et al. [50]. More precisely, our two main points of interest were the sharing decisions of the creators and the allocation rule of the platform, which are novel aspects of our paper. The adapted result is as follows:
Adapted Theorem. Assume power cost functions $c_{i}(x_{i})=a_{i}x_{i}^{\theta}$ for all $i$. Fix any $\rho\in[0,1]$ and let $\mu^{\prime}=\mu(1+\alpha\rho)$. For any sufficiently large $n$ and $\mu$, $\left\|\mathcal{X}(\rho)\right\|$ satisfies:
$\frac{C_{\theta}}{2\alpha+2}<\frac{(\left\|\mathcal{X}(\rho)\right\|)^{\theta-\gamma}}{\mu^{\prime}\cdot\left\|\mathbf{a}^{-1}\right\|_{\frac{1}{\theta-1}}}<C_{\theta},$
where $C_{\theta}$ is a constant depending only on $\theta$, $\mathbf{a}^{-1}=(a_{1}^{-1},\dots,a_{n}^{-1})$, and $\left\|\mathbf{a}^{-1}\right\|_{\frac{1}{\theta-1}}=\left(\sum_{i=1}^{n}(a_{i}^{-1})^{\frac{1}{\theta-1}}\right)^{\theta-1}$.
We now proceed to derive the approximation. We rely on the key relationships between $\left\|\mathcal{X}(\rho)\right\|$ and its lower bound $X_{LB}(\rho)$ established in the theorem above, which guarantees that for any $\rho\in[0,1]$:
$X_{LB}(\rho)\leq\left\|\mathcal{X}(\rho)\right\|\leq R\cdot X_{LB}(\rho),$
where $R=(2\alpha+2)^{1/(\theta-\gamma)}$ and $X_{LB}(\rho)=\left(\frac{C_{\theta}\mu(1+\alpha\rho)\left\|\mathbf{a}^{-1}\right\|_{\frac{1}{\theta-1}}}{2\alpha+2}\right)^{\frac{1}{\theta-\gamma}}$. Since the platform’s revenue is $U_{p}(\mathcal{X}(\rho),\mathbf{1};f_{\rho})=\frac{\mu\alpha}{1+\alpha}(1-\rho)(\mathcal{X}(\rho))^{\gamma}$, these quality bounds imply corresponding revenue bounds:
$V_{LB}(\rho)\leq U_{p}(\mathcal{X}(\rho),\mathbf{1};f_{\rho})\leq R^{\gamma}\cdot V_{LB}(\rho),$
@@
$V_{LB}(\rho)$ $=\frac{\mu\alpha}{1+\alpha}(1-\rho)\left[\left(\frac{C_{\theta}\mu(1+\alpha\rho)\left\|\mathbf{a}^{-1}\right\|_{\frac{1}{\theta-1}}}{2\alpha+2}\right)^{\frac{1}{\theta-\gamma}}\right]^{\gamma}$